# Reva Schwartz1    Carina Westling2    Morgan Briggs3    Marzieh Fadaee4    Isar Nejadgholi5    Matthew Holmes6    Fariza Rashid7    Maya Carlyle8    Afaf Taïk9    Kyra Wilson10    Peter Douglas11    Theodora Skeadas12    Gabriella Waters1,13    Rumman Chowdhury14    Thiago Lacerda3

### 03 Mar 2026 v2

@@ 2.1.3 Stage 3: Evaluation execution
In Stage 3, testing is carried out through real‑world trials of AI systems with large, diverse participant groups so that findings can generalize across user categories and contexts [19]. It also sets clear plans for who is eligible to participate and how samples are drawn, in line with evaluation requirements and human subjects research protocols and with attention to participant incentives and expected attrition. AI’s higher‑order, ecosystem‑level effects influence both users and non‑users of AI systems. Sampling can therefore be designed to include people who do not use AI, for example through stratified recruitment and oversampling, so their behaviors are well represented. Their participation can then be up‑weighted in analysis to counter adopter‑centric biases in existing data and improve generalizability to groups currently underserved by chat models [69, 33].
For example, achieving this in EdTech might require enrolling teachers from different schools, with different seniority levels/ experience with AI [51].
Figure 6: Figure 6. Stage 3 of the lifecycle in the EdTech example: executing real‑world trials to observe the construct in practice.
Participants are trained prior to testing to make sure they can meaningfully and safely exercise the system under evaluation, understand their role in the study, and are aware of the system’s intended use and limitations. A shared code of conduct is used to inform participants of applicable ethical considerations, data collection practices, acceptable use, and how to report or escalate any harms or unexpected behavior they may experience [46]. The code of conduct also includes plain‑language descriptions of core functionality, known failure modes, and key limitations. Depending on the evaluation design, testers may also be trained on how to conduct adversarial testing.
Fully automated tests can be run alongside real-world trials using an identical test harness, enabling consistent and directly comparable evaluation outcomes and helping to reveal the contextual and higher‑order effects that are most consequential for stakeholders. All decisions across the lifecycle framework are tracked and documented so that evaluation outcomes can be interpreted based on testing mode, who participated and under what conditions. The hypotheses, metrics, sampling plans, and protocols defined in the design stage often need adjustments due to changes that occur in the execution stage. Typical adaptations in test execution include recruitment shortfalls with less people showing up, and technical difficulties leading to altered usage patterns or missing data. Documenting such deviations is necessary for a sound analysis in the subsequent steps or determining whether experiments need to be repeated or newer ones added.
@@ 2.1.4 Stage 4: Analysis
Stage 4 begins with scenario design, which creates the scaffolding needed to make the constructs of interest observable and measurable [54, 68]. Each construct is translated into a high-level description of in‑context behaviors that can be expected to realistically occur in the deployment environment. For example, systematically observing whether a teacher reviews AI-generated quizzes for required learning outcomes allows for concrete analysis. This behavior can be measured using quiz outputs, student chatbot transcripts, and documented follow-up actions, providing data to assess both short-term and long-term effects across classrooms. Proxy scenarios can also be used to evaluate hard‑to‑observe higher‑order effects, such as over‑reliance on chatbot engagements, without directly exposing participants to harmful content or experiences.
Scenarios also anchor the development of other analysis tools, including scoring rubrics, scoring criteria, and data markup methods. Scoring criteria are designed to connect the elicited evaluation behaviors to what actually materializes in context. Rubrics are used to consistently map raw observations onto outcome metrics that are tied to stakeholder concerns. Annotation and markup of output data can be fully automated, fully manual, or participant-driven but must use universally understood definitions for the observed behaviors to be sure that everyone applies and interprets the definitions consistently throughout the markup process. Figure 7 highlights Stage 4, translation of constructs into observable behaviors and indicators.
Figure 7: Figure 7. Stage 4 of the lifecycle in the EdTech example: translating the construct into observable teacher behaviors and classroom signals.
This stage combines qualitative real‑world results from classroom trials and large‑scale model runs so that each informs the other in context rather than standing alone. Observed behaviors and coded patterns from contextualized student-chatbot transcripts and other collected data are compared to automated metrics from scenario‑matched test suites. Because both evaluation modes draw on the same underlying constructs, the combined output measures support contextual awareness by revealing not only raw system outcomes, but the conditions under which model behavior is likely to contribute to over‑reliance, cognitive offloading, or other secondary effects at scale. This contextually enriched information supports broader inferences about tertiary effects that link concrete classroom practices to longer‑term shifts in learning outcomes, curricula, and teaching effectiveness.
@@ 2.1.5 Stage 5: Insights
Interpretation of data re-contextualizes and extrapolates the findings that are produced by analysis. By using matched scenarios in the education example, the data from local tests can be aggregated with large-scale model runs to enable longitudinal projection and interrogation of tertiary effects. The interpretation process can also engage participants and stakeholders in guided exploration of the data produced in local tests to re-contextualize findings and formulate insights for dissemination to relevant stakeholders and communities.
Figure 8: Figure 8. Stage 5 of the lifecycle in the EdTech example: using observed system outputs to generate insights that inform subsequent teacher actions.
The choice of method is context-sensitive and should consider audiences, reach and impact. Concepts of interest identified in Stage 1 can continue to guide interpretation for insight generation. Where possible, continuous stakeholder engagement, from context specification through to insight generation and dissemination, enhances consistency and impactful insights.
The interpreted findings and insights can be formulated and presented to optimize reach and impact for intended audiences and platforms or channels. Figure 8 highlights how Stage 5 uses system outputs to generate insights and inform actions in the EdTech example. Care should be taken to reduce the use of highly technical language to ensure that evaluation outcomes can be easily leveraged by all relevant stakeholders. Stakeholders can assist in identifying the type of platforms and publications that are a match in community type and publication tone so materials can be optimized for reach and impact. Typically, this might include industry periodicals, newsletters, and other forms of internal and/or community-facing publications, conferences, peer-reviewed journals (academic and non-academic), and news media. It can also include formal and informal networks and community groups and organizations. With the goal of supporting stakeholders beyond the stack, insights should conform to the type and style of communication that is established within the organization or community in question. Other considerations include how to record, format, store, and disseminate findings.
@@ 2.1.6 Stage 6: Continuous monitoring
Practically, effective monitoring in education is a hybrid process[59]. Automated systems can collect signals such as usage patterns and learning outcomes. Human and institutional review is required to interpret these signals in relation to local goals, constraints, and unintended consequences, and to determine whether corrective action, redesign, or policy updates are warranted. The goal is a recurring cycle—observe, interpret, diagnose, act—that supports continuous improvement without collapsing monitoring into continuous surveillance of individuals or real-time behavioral control [39]. Accordingly, monitoring follows principles of data minimization, proportionality, and purpose limitation [10, 45], , ensuring that oversight remains legitimate, auditable, and aligned with institutional governance and privacy obligations. Findings from continuous monitoring feed back into renewed context specification and evaluation design, allowing constructs, scenarios, and metrics to be revised as deployment conditions and stakeholder priorities change.
@@ 3 Discussion and Future Work
CIRCLE reframes AI evaluation as a construct-centered, lifecycle process that links stakeholder-defined concerns to contextualized scenarios, mixed-methods testing, and continuous monitoring. Rather than treating evaluation methods such as benchmarks, red teaming, and field trials as isolated activities, the framework integrates them into a single pipeline that ties every metric to a named construct, observable behavior, and stakeholder-relevant outcome. This structure makes it possible to collect observational data about the concepts that matter to stakeholders outside the AI stack–such as secondary and tertiary effects–, and to generate evidence that informs questions about whether and how systems should be adopted and governed in their own settings.
@@
Implementing the CIRCLE lifecycle introduces trade-offs regarding cost and complexity. Unlike automated benchmarks, collecting context-aware information requires organizational maturity, interdisciplinary expertise, and time-intensive human-subjects protocols, and is inherently slower and more resource‑intensive than running in silico test scripts. At the same time, these investments directly support stronger forms of measurement validity that are difficult to achieve with purely model-centric evaluations.
CIRCLE enhances ecological validity by structuring evaluations in settings that approximate everyday life (e.g., classrooms) rather than in tightly controlled, artificial test environments. It strengthens construct validity by explicitly eliciting stakeholder concerns in Stage 1 and naming them as constructs in Stages 2 and 4 (e.g., over-reliance, cognitive offloading, displacement in classrooms), then designing scenarios, coding schemes, and metrics that instantiate these constructs within actual workflows.
@@
CIRCLE also supports criterion validity by aligning outcome measures with real-world stakeholder outcomes such as durable ROI, shifts in workload, or changes in error profiles, making evaluation metrics more reliable predictors of deployment performance. Consequential validity is improved by systematically tracking benefits and harms through real-world interactions across the lifecycle, rather than inferring them only from theory or in silico testing results. Finally, internal validity is strengthened through deliberate design choices in Stage 2, such as comparison groups and quasi-experimental designs, and through careful documentation of protocols and deviations during execution, which help minimize alternative explanations and experimental confounds.
@@
Future efforts will need to reduce the friction of capturing contextual detail at scale, for example by developing validated user simulators and other tooling grounded in empirical lifecycle data. Adoption of contextually aware insights about higher-order effects also requires a shift in institutional incentives. Regulators will need to place value on deployment-level evidence in addition to model documentation. This may lead to enterprise investment in comprehensive measurement infrastructure that tracks value rather than just usage, and shared libraries of operationalized constructs (e.g., cognitive offloading) to make these evaluations interoperable. Without this shift, the ecosystem will remain stuck measuring abstract capabilities while missing the materialized risks of deployment.

### 18 Mar 2026 v3

@@ 2.1.3 Stage 3: Evaluation execution
Testing is carried out in Stage 3, through real‑world trials of AI systems with large, diverse participant groups so that findings can generalize across user categories and contexts [19] (illustrated in Figure 6). This stage also defines participant sampling based on evaluation requirements and in line with human subjects research protocols, participant incentives, and expected attrition. Since AI’s higher‑order effects can impact both users and non‑users, evaluators can use stratified sampling and oversampling to recruit people who do not use AI so their behaviors are represented. Non‑user responses can then be up‑weighted in analysis to reduce adopter‑centric bias and better represent groups currently underserved by chat models [68, 32]. For example, in EdTech this might require enrolling teachers from different schools and with different seniority levels and experience with AI [50].
Figure 6: Stage 3 of the lifecycle in the EdTech example: executing real‑world trials to observe the construct in practice.
Participants receive training before testing so they can use the system safely and understand their role and the system’s intended use. They are provided informed consent and a code of conduct covering ethical considerations, data collection, acceptable use, and how to report or escalate harms or unexpected behavior [45]. Codes of conduct should also describe core system functionality, known failure modes, and key limitations in plain language. Depending on the evaluation design, participants may also be trained to conduct adversarial testing.
Fully automated tests can run alongside participant trials using the same test harness, enabling directly comparable evaluation outcomes and providing a more comprehensive view of the deployment environment. Decisions across the lifecycle should be tracked and documented so evaluation outcomes can be interpreted in light of testing mode, conditions, and participant category. Hypotheses, metrics, sampling plans, and protocols defined in the design stage typically require adjustment due to factors such as recruitment shortfalls or technical difficulties that lead to unexpected usage patterns or missing data. Documenting these deviations is essential for sound analysis in subsequent stages and for deciding whether to repeat, adapt, or add experiments.
@@ 2.1.4 Stage 4: Analysis
Stage 4 supports scenario design by creating the scaffolding needed to make constructs of interest observable and measurable [53, 67]. For example, systematically observing how often a teacher reviews AI-generated quizzes for required learning outcomes makes the over-reliance construct concrete and analyzable. This behavior can be tracked through student quiz results, chatbot transcripts, and records of any follow‑on actions–yielding insights into higher-order effects within and across classrooms. Proxy scenarios can be used to assess hard‑to‑observe higher‑order effects without directly exposing participants to harmful content or experiences.
Scenarios anchor the development of other analysis tools, including scoring rubrics, scoring criteria, and data markup methods. Scoring criteria are designed to connect the elicited evaluation behaviors to what actually materializes in context. Rubrics are used to consistently map raw observations onto outcome metrics–all of which are tied to stakeholder-articulated priorities. Annotation and markup of output data can be fully automated, fully manual, or participant-driven, but must use universally understood definitions for the observed behaviors to be sure that everyone applies and interprets the definitions consistently throughout the markup process. Figure 7 highlights Stage 4, translation of constructs into observable behaviors and indicators.
Figure 7: Stage 4 of the lifecycle in the EdTech example: translating the construct into observable teacher behaviors and classroom signals.
Analysis can draw on qualitative results from classroom trials and large‑scale model runs so each informs the other in context rather than standing alone. Observed behaviors and coded patterns from contextualized student‑chatbot transcripts and other data are then compared to automated metrics from scenario‑matched test suites. Because both evaluation modes draw on the same constructs, their combined measures can connect specific deployment conditions, model behavior under those conditions, and the secondary effects that emerge at scale. This context‑rich information, in turn, supports broader inferences about tertiary effects that link concrete classroom practices to longer‑term shifts in learning outcomes, curricula, and teaching effectiveness.
Definitions for key testing concepts described in Stage 4 are listed in the “Testing Scenario Concepts” box below.
@@ 2.1.5 Stage 5: Insights
Deriving evaluation insights involves putting analysis findings back into context and extrapolating them to likely future conditions. Depending on how the participant and automated runs are designed, their combination can support longitudinal projections and examination of tertiary effects. Bringing participants and stakeholders into guided discussions of local test results helps re‑contextualize the findings and translate them into insights that meet their interpretive needs.
Figure 8: Stage 5 of the lifecycle in the EdTech example: using observed system outputs to generate insights that inform subsequent teacher actions.
Interpretive methods are context-sensitive and should consider audiences, reach, and impact and can make use of the oncepts identified by stakeholders in Stage 1. Findings should be presented in ways that match the needs and communication styles of the intended audiences, relying on plain language for maximum uptake beyond technical audiences. Figure 8 shows how EdTech evaluation outputs can be used in Stage 5 to generate insights and inform the stakeholder audience. Stakeholders can support selection of appropriate channels—such as industry periodicals, newsletters, internal reports, conferences, journals, news media, and community networks—and shape tone and format so materials are easy to find, understand, and act on.
@@ 2.1.6 Stage 6: Continuous monitoring
Practically, effective monitoring in education is a hybrid process[58]. Automated systems can collect signals such as usage patterns and learning outcomes. Human and institutional review is required to interpret these signals in relation to local goals, constraints, and unintended consequences, and to determine whether corrective action, redesign, or policy updates are warranted. The goal is a recurring cycle—observe, interpret, diagnose, act—that supports continuous improvement without collapsing monitoring into continuous surveillance of individuals or real-time behavioral control [38]. Accordingly, monitoring follows principles of data minimization, proportionality, and purpose limitation [10, 44], ensuring that oversight remains legitimate, auditable, and aligned with institutional governance and privacy obligations. Findings from continuous monitoring feed back into renewed context specification and evaluation design, allowing constructs, scenarios, and metrics to be revised as deployment conditions and stakeholder priorities change.
@@ 3 Discussion and Future Work
CIRCLE treats AI evaluation as a construct‑centered, lifecycle process that links stakeholder priorities to contextualized scenarios, mixed‑methods testing, and ongoing monitoring. Instead of treating benchmarks, red teaming, and field trials as separate activities, it integrates them into one pipeline that ties every metric to a named construct, an observable behavior, and a stakeholder‑relevant outcome. This makes it possible to generate evidence about the concepts that matter to people outside the AI stack so they can make informed decisions about whether and how systems should be adopted and governed in their own settings.
@@
Implementing the CIRCLE lifecycle introduces trade-offs regarding cost and complexity. Unlike automated benchmarks, collecting context-aware information requires organizational maturity, interdisciplinary expertise, and time-intensive human-subjects protocols, and is inherently slower and more resource‑intensive than running in silico test scripts. These investments, however, strengthen measurement validity and enable organizations to surface higher‑order, context‑specific insights about downstream and tertiary effects that directly inform deployment, governance, and decommissioning decisions.
@@
CIRCLE enhances ecological validity by structuring evaluations in settings that approximate everyday life rather than in tightly controlled, artificial test environments. It strengthens construct validity by explicitly eliciting stakeholder concerns in Stage 1,naming them as constructs in Stages 2 and 4, then designing scenarios, coding schemes, and metrics that instantiate them for evaluation. Criterion validity is supported by aligning evaluation outcome measures with real‑world outcomes of interest, enabling more reliable predictors of deployment performance. Consequential validity is improved by systematically tracking benefits and harms through real-world interactions across the lifecycle, rather than inferring them only from theory or in silico testing results. Finally, internal validity is strengthened through deliberate design choices in Stage 2, such as comparison groups and quasi-experimental designs, and through careful documentation of protocols and deviations during execution, which help minimize alternative explanations and experimental confounds.
@@
Future efforts will need to reduce the friction of capturing contextual detail at scale, for example by developing validated user simulators and other tooling grounded in empirical lifecycle data. Adoption of context‑sensitive insights about higher‑order effects will also require shifts in institutional incentives so that deployment‑level evidence is valued alongside model documentation. Investing in measurement infrastructure, including shared libraries of operationalized constructs, can help make these evaluations more interoperable across settings.
4 Conclusion
This paper presented CIRCLE as a construct‑centered lifecycle framework for evaluating AI systems from a real‑world perspective to address the gap between static, model‑centric benchmarks and dynamic deployment outcomes. The six stages coordinate activities such as stakeholder elicitation, evaluation design, real‑world testing, analysis, and monitoring into a traceable pipeline that produces systematic knowledge about how AI systems actually function in practice. This enables decision‑makers beyond the AI development stack to ground adoption and governance decisions in evidence about materialized effects, not just theoretical capabilities.